# Structural Change, Employment, and Inequality in Europe: an Economic Complexity Approach

### v1

@@ Appendix A Data reconstruction
The Eurostat’s Structural Business Survey data used to compute the employment based EFC measures used in this paper are one of the most comprehensive and detailed data sources of employment in Eurpean industries, classified at a fine level of disaggregation (NACE 4 digit). However, the dataset presents a large number of missing values, which hamper the construction of a specialisation matrix as explained in Section 4.1. In fact, the BiWCM and EFC algorithm require both a matrix with complete information.
In order to address this shortcoming in the SBS data, we have pursued, compared and validated seven different interpolation strategies, in order to retrieve a good approximation of the missing data. The first step has been to attempt the reconstruction of the SBS data following 7 alternative strategies, applied to the country-industry time series having a maximum 7 missing data points [^5] In the case of fully missing series, the country-industry couple has been excluded. In case of 8 out of 9 missing data points, the value has been assumed to be constant across the time series, extrapolating at constant values from the only available data point in the series.:
@@
  • Reflect frequency in original data ± 5%,
  • Reflect frequency in original data + 10%
@@
In order to pick the data reconstruction strategy that best approximates the observed data, we compare the Mean Absolute Error of each strategy across the 5 validation samples. As shown by table 7, the strategy that minimises the prediction error across the different validation subsamples is the first – i.e. the one that interpolates internal values linearly, and extrapolates external values taking the first and last available data point as constant, both backward and forward.
Table 7: Mean Absolute Errors

### v2

@@ Appendix A Data reconstruction
The Eurostat’s Structural Business Survey data used to compute the employment based EFC measures used in this paper are one of the most comprehensive and detailed data sources of employment in European industries, classified at a fine level of disaggregation (NACE 4 digit). However, the dataset presents a large number of missing values, which prevent the construction of a specialisation matrix as explained in Section 4.1. In fact, the BiWCM and EFC algorithm require both a matrix with complete information.
In order to address this shortcoming in the Structural Business Survey data described in Section 3, we have pursued, compared, and validated seven different interpolation strategies, in order to retrieve a good approximation of the missing data. We have attempted the reconstruction of the SBS data following seven alternative strategies, applied to the country-industry time series having a maximum seven missing data points. [^8] In the case of fully missing series, the country-industry couple has been excluded. In case of eight out of nine missing data points, the value has been assumed to be constant across the time series, extrapolating at constant values from the only available data point in the series. Below we summarise the seven distinct strategies:
@@
  • Reflect frequency in original data $\pm$ 5%,
  • Reflect frequency in original data $\pm$ 10%,
@@
Table A1: Mean Absolute Errors
@@
  • Notes: The table reports mean absolute errors (MAE) across different levels of missing data (NA). Bold values indicate the lowest MAE within each subsample.
In order to choose the data reconstruction strategy that best approximates the observed data, we compare the Mean Absolute Error of each strategy across the 5 validation samples. As shown by Table A1, the strategy that minimises the prediction error across the different validation subsamples is strategy number 1. Therefore, the final employment dataset used to construct the specialisation matrix interpolates internal missing values linearly, and extrapolates external values taking the first and last available data point as constant, both backward and forward. According to our test, this strategy is the one that is most likely replicate the real data.
Appendix B Methodological notes
B.1 Differences between Revealed and Inferred Comparative Advantage
The primary rationale for utilising a Comparative Advantage index is to establish a measure that isn’t contingent on the relative strength of diverse countries and industries, which are inherently incomparable in terms of size. Revealed Comparative Advantage (RCA) and Inferred Comparative Advantage (ICA) share the same conceptual core: both assess whether the observed extensive quantity, a volume, $V_{ci}$ of actor $c$ in activity $i$ exceeds an expected value under a benchmark. In this work, the volume is the labour force, the actor is a country, and the activity is an industry (NACE 4 digits). In general terms, a comparative advantage index can be cast as an observed-to-expected ratio:
$\mathrm{CA}_{ci}=\frac{V_{ci}}{\mathbb{E}[V_{ci}]}\,$ (12)
Where observed term is the share of country $c$ in industry $i$, and the expected term is – in general terms – an expectation of volume in industry $i$ based on the whole sample of countries. The difference between RCA rests in the way in which the expectation $\mathbb{E}[V_{cp}]$ is constructed – namely on the constraints imposed on the expected value of volume.
The traditional RCA measure proposed by Balassa (1965), specifically, contrasts the actual volume of a country with an anticipated value derived under the assumption of equalising country and sector sizes (Vollrath, 1991; Krantz et al., 2018). In other words, the share of country $c$ in industry $i$ is compared to the average share of all countries in the same industry $i$:
$\mathrm{RCA}_{ci}=\frac{V_{ci}/\sum_{i^{\prime}}V_{ci^{\prime}}}{\left(\sum_{c^{\prime}}V_{c^{\prime}i}\right)/\sum_{c^{\prime}i^{\prime}}V_{c^{\prime}i^{\prime}}}\,$, (13)
which is equivalent to comparing $V_{ci}$ to a factorised expectation based on marginal totals: let $V_{c}=\sum_{i}V_{cp}$, $V_{p}=\sum_{c}V_{ci}$, and $V=\sum_{c,i}V_{ci}$. Then
$\mathrm{RCA}_{ci}=\frac{V_{ci}}{\mathbb{E}_{\mathrm{RCA}}[V_{ci}]}\qquad\text{with}\qquad\mathbb{E}_{\mathrm{RCA}}[V_{ci}]=\frac{V_{c}\,V_{i}}{V}\,.$ (14)
Therefore, RCA implicitly adopts an independence baseline where the expected intensity in $(c,i)$ is determined solely by the overall scale of $c$ and the overall prevalence of $i$. This assumption is generally considered reasonable when the volumes are of similar magnitude and no correlation is present in the system.
ICA (Bruno et al., 2023) avoids this naive assumption, which may not apply to all types of data. ICA retains the same observed-to-expected structure but replaces the implicit baseline with an explicit probabilistic null model derived by maximum entropy. Specifically, the expectation $\mathbb{E}[V_{ci}]$ is computed under the Bipartite Weighted Configuration Model (BiWCM), i.e. the maximum-entropy ensemble of weighted bipartite networks that is maximally random subject to reproducing the total marginal volumes as constraints (i.e. the node strengths, a.k.a. the row and column sums) in expectation. Denoting the BiWCM-implied expected weight by $\mathbb{E}_{\mathrm{BiWCM}}[V_{ci}]$, ICA is defined as
$\mathrm{ICA}_{ci}=\frac{V_{ci}}{\mathbb{E}_{\mathrm{BiWCM}}[V_{ci}]}\,.$ (15)
The key methodological analogy is that both RCA and ICA quantify comparative advantage as a deviation from an expected value anchored to aggregate information. The key methodological difference is that RCA uses a closed-form expectation $\mathbb{E}_{\mathrm{RCA}}[V_{ci}]=V_{c}V_{i}/V$ implied by the volume independence assumption, whereas ICA uses the maximum-entropy expectation $\mathbb{E}_{\mathrm{BiWCM}}[V_{ci}]$, obtained from an explicit entropy-maximising (equivalently, maximum-likelihood) null model consistent with the imposed constraints. This approach naturally supports uncertainty quantification and statistical validation when needed, while minimising statistical bias.
Finally, the comparative advantage index is binarised whether it is larger or smaller of a threshold. The value is naturally set to 1, stating that we assume a comparative advantage whenever a country in a product does more than expected by the null model. Therefore we have:
$M_{c,i}=\begin{cases}1&\mathrm{ICA}_{c,i}\geq 1\\ 0&\text{otherwise}\end{cases}$ (16)
Economically, ICA can be interpreted as asking whether a country employs more labour in a given industry than would be expected purely from its overall size and the global importance of that industry, once the structural constraints of the entire country–industry system are taken into account. In this sense, ICA identifies industries where the observed allocation of labour cannot be explained by aggregate scale effects alone, but instead reflects a genuine concentration of productive capabilities or specialisation relative to the structure of the global economy.
Figure A1 compares the diversification rankings obtained using the ICA and RCA frameworks. The results reveal substantial differences between the two measures: for instance, countries such as the UK and Spain display almost opposite levels of relevance depending on the metric adopted. In contrast, Poland, Hungary, the Czech Republic, and Bulgaria rank at the top under the RCA framework and also remain among the highest‑ranked countries when evaluated with ICA. These discrepancies stem from the uneven distribution of the labour force across sectors and countries. Such imbalances influence how each CA framework captures diversification and the underlying capabilities. Indeed, when examining the distribution of diversification directly, it becomes evident that the RCA tends to produce a more uniform and less differentiated picture, whereas the ICA preserves stronger heterogeneity across countries.
Figure A1: Comparison between ICA and RCA diversification ranking in 2014.
  • Note: The two horizontal bar charts display each country’s level of diversification as measured by the two comparative advantage indices.
B.2 Making Fitness comparable over time
The EFC algorithm computes a fixed point to determine the actual values of the Fitness and Complexity indices for each year independently. While the rankings of the indexes have a clear interpretation, their numerical values are relative to a scale that in unknown in the original formulation of the algorithm. What is the meaning of a Fitness $F_{c}=5$? That depends on the scale we are using to measure the index. Such scale, if not fixed by the user, depends on the density of the input matrix and on the distribution of the points among rows and columns. Such quantities change over time and, with them, the reference scale of the values changes as well. Consequently, in principle, it is not possible to compare the index values across different years due to the dynamic changes in the bipartite networks. Negative changes of a country’s Fitness may not reflect an actual loss of capabilities or competitiveness but a change in $M_{c,i}$ such as other countries gaining new products and changing the total density of the matrix.
To address this issue, we employ a technique, the "dummy trick", to fix a constant reference scale for the numerical values as elaborated in (Mazzilli et al., 2024). We add to $M_{c,i}$ an additional row of all ones (the "dummy" row) which represents a fictitious country specialised in all industries. We rescale the Fitness values dividing all of them by the Fitness of the ’ideal’ country, that is now fixed to $1$. Thus, in this version of the algorithm all $F_{c}<1$ for all real countries. The scale of reference is fixed by the ’ideal’ country that would export every product, and Fitness values indicate how close each country is to it. This approach establishes a fixed reference scale over time, enabling the comparison of Fitness and Complexity values across different years.